OpenStudy (anonymous):

Find the slope of the line passing through (11, -2) and (2, -2).

OpenStudy (anonymous):

use the slope formula \[\frac{ y^2-y^2 }{ x^2-x^1 }\] \[\frac{ 2-(-2) }{ 2-11 }\] =\[\frac{ ?? }{ ?? }\]